If 2.39 mol of ethane undergoes combustion how many moles of O2 are required answer in units of mol


1
Expert's answer
2021-02-26T06:02:52-0500

The reaction of ethane combustion: is as following:

2C2H6 + 7O2 = 4CO2 + 6H2O

Since the number of moles of ethane and oxygen must be equal, the number of moles of oxygen:

n(O2) = (n(C2H6) × 7) / 2 = (2.39 mol × 7) / 2 = 8.365 mol


Answer: 8.365 mol